OpenStudy (anonymous):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 3,250n - 9,000 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (anonymous):

PLEASE HELP ILL MEDAL

OpenStudy (anonymous):

by some miracle this factors as \[-250 (n-4) (n-9)\] so the zeros should be easy to find, solve \[(n-4)(n-9)=0\]

OpenStudy (anonymous):

I already have Part A.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

oh find the first coordinate of the max via \(x=-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case it is \(\frac{13}{2}\) or \(6.5\)

OpenStudy (anonymous):

So thats Part B?

OpenStudy (anonymous):

not coincidentally half way between the zeros

OpenStudy (anonymous):

to find the second coordinate of the max, evaluate the function at \(\frac{13}{2}\) let me know what you get then i can show you what it looks like when the square is completed

OpenStudy (anonymous):

where would you plug it in?

OpenStudy (anonymous):

here \[P(n) = -250n^2 + 3,250n - 9,000\] use a calculator

OpenStudy (anonymous):

P(n)=-250(n-4)(n-9)

OpenStudy (anonymous):

what is \[P(\frac{13}{2})\]?

OpenStudy (anonymous):

\[\frac{ 13P }{ }\]

OpenStudy (anonymous):

over 2 sorry

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i sense you are lost \[P(n)\] is a function \[P(\frac{13}{2})\] is a number since \[P(n) = -250n^2 + 3,250n - 9,000\] then \[P(\frac{13}{2}= -250(\frac{13}{2})^2 + 3,250(\frac{13}{2}) - 9,000\]

OpenStudy (anonymous):

oh okay, so solve the last equation?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

solve in this case means compute it use a calculator

OpenStudy (anonymous):

here i will show you an easy on line one to use

OpenStudy (anonymous):

so, that would be Part B?

OpenStudy (anonymous):

that means is "vertex form" you have \[P(n)=-250(n-6.5)^2+1562.5\]

OpenStudy (anonymous):

i resorted to decimals because the fractions were getting annoying and yes, that is the answer to B

OpenStudy (anonymous):

okay, what about Part C?

OpenStudy (anonymous):

Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) the maximum is \(1562.5\) which occurs if \(n=6.5\) and therefore the vertex form is \[P(n)=-250(n-6.5)^2+1562.5\] you may notice that i did not actually complete the square, it is easier to find the vertex without doing that

OpenStudy (anonymous):

Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (anonymous):

already found it, it is the first coordinate of the vertex written as a line \[x=\frac{13}{2}\] or \[x=6.5\]

OpenStudy (anonymous):

13/2?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Can you help me on another?

OpenStudy (anonymous):

sure if it is quick

OpenStudy (anonymous):

Kim is studying the sale of a particular brand of cereals from the year 2000 to 2010. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 7t + 69 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points)

OpenStudy (anonymous):

I already have Part A & B. Just need Part C

OpenStudy (anonymous):

\[S(t) = t^2 + 7t + 69\] the y intercept is 69, the millions of dollars when you start, in 2000

OpenStudy (anonymous):

oh you got that ok

OpenStudy (anonymous):

Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points) this always confuses me i assume the first year is \(2000\) so the fourth year is \(2003\) then the average rate of change is the slope \[\frac{s(3)-s(0)}{3-0}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

\[S(t) = t^2 + 7t + 69\] \[S(0)=69\] \[S(3)=99\] \[\frac{S(3)-S(0)}{3}=\frac{99-69}{3}=\frac{30}{3}=10\]

OpenStudy (anonymous):

yw